# GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks

### v1

@@ Abstract
In this paper, we establish a benchmark for evaluating large language models (LLMs) on multi-step geospatial tasks relevant to commercial GIS practitioners. We assess seven leading commercial LLMs (Sonnet 3.5 and 3.7, Haiku 3.5, Gemini 2.0, GPT-4o, GPT-4o mini, and o3-mini) using a simple tool-calling agent equipped with 23 geospatial functions. Our benchmark comprises tasks across four categories of increasing complexity, with both solvable and intentionally unsolvable tasks to test hallucination rejection. We develop an LLM-as-Judge evaluation framework to compare agent solutions against reference implementations. Results show Sonnet 3.5 and GPT-4o achieve the best overall performance, with Claude models excelling on solvable tasks while OpenAI models better identify unsolvable scenarios. We observe significant differences in token usage, with Anthropic models consuming substantially more tokens than competitors. Common errors include misunderstanding geometrical relationships, relying on outdated knowledge, and inefficient data manipulation. The resulting benchmark set, evaluation framework, and data generation pipeline are released as open-source resources, providing one more standardized method for ongoing evaluation of LLMs for GeoAI.
GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks
Varvara Krechetova World Bank New York, USA                        Denis Kochedykov J.P.Morgan Chase New York, USA
1 Introduction
@@
In this paper, we create a benchmark set and describe a simple LLM-as-Judge based methodology to evaluate the performance of several leading commercial large language models (LLMs) with function-calling abilities for solving a range of multi-step geospatial tasks. Geospatial tasks are distinct in that they require LLMs to have good factual geographical knowledge, understanding how to work with spatially defined data, georeferenced data, what a coordinate reference system is, understanding what are various types of spatial or geodata. We used a simple tool-calling agent to specifically assess the LLMs capabilities and have tested several models: Sonnet 3.5 and 3.7, Haiku 3.5, Gemini 2.0, GPT-4o, GPT-4o mini, o3-mini.
@@
Our focus is on the commercial GIS practitioner requirements, which influence the choice of tasks, models and evaluation framework. Unlike other adjacent benchmarks, we focus on tasks that are closer to what GIS practitioners usually need for socio-economic analysis and mapping for reports, news articles, memos, and the like. Practitioners would use LLM equipped with GIS tools in a chat environment to get series of maps, numbers, and lists of locations. The tasks in the benchmark set reflect this – asking to make various types of map, to find geographical objects by some conditions, count objects within certain geographic boundaries, etc. Also, importantly for practitioners, we test the ability of LLMs to correctly recognize when the task cannot be solved with given tools and inform the user. This reduces the risk of hallucinated outputs. Next, we benchmark most current advanced commercial LLMs used in industry setting. One requirement of such industry usage is that the variability of outputs is not desired, so the models were tested at temperature 0. Lastly, the evaluation method we propose allows one to automate the ongoing evaluation which is increasingly important in practice with frequent models updates in commercial settings.
The benchmark set and code are publicly released at https://github.com/Solirinai/GeoBenchX.
@@
The paper is structured as follows. First, we describe the task-solving agent and the set of tools/datasets. Second, we describe how the benchmark set of tasks was created and its composition. Third, we detail the approach to evaluation of the results of the task-solver agent using an LLM-as-Judge evaluator agent and testing the evaluator performance. Lastly, we compare the performance metrics of the solver agent based on different LLMs and the cost efficiency in terms of token usage.
2 Related works
@@
Research in the field of GeoAI, a part of spatial data science at the intersection with Artificial Intelligence (AI) Janowicz et al. (2020), got a boost with recent advances in generative AI and the deployment of accessible LLMs. GeoAI researchers actively explore the capabilities of the available models in relation to geospatial data science, design their applications, and methods to improve their performance. The avenues explored include analysis of the ability of LLM (GPT-4) for spatial reasoning in travel networks, distance and elevation estimations, route and supply chain analysis Roberts et al. (2023); creating autonomous LLM-based GIS applications (LLM-Geo) Li and and (2023), GIS Copilot for GIS software Akinboyewa et al. (2024), ability LLMs to correctly understand and apply GIS toolsZhang et al. (2024), integrating LLM and cartography tools into MapGPT He et al. (2024), combining several LLMs to increase the efficiency of spatial data processing in the multi-agent framework for automated processing of vector data in shapefile format Lin et al. (2024), benefits of generative AI agents in smart mobilityXu et al. (2024), fine-tuning LLM to process natural language queries into executable codeMansourian and Oucheikh (2024). There is a benchmarking of LLM performance in spatial tasks in progress, with analysis of LLMs’ geospatial knowledgeBhandari et al. (2023); Renshaw et al. (2025), of GIS coding abilitiesGramacki et al. (2024), the recent study testing models on several groups of tasks that require geographic, GIS, and spatial knowledge, including changes in their performance in response to changes in prompts Liuchang Xu et al. (2025).
3 Agent, tools, and datasets
@@
The experiment setup is based on the pre-built Langgraph ReAct agent LangChain AI (2024) equipped to call one of the selected LLMs with a list of attached tools. We intentionally choose a simple agent architecture to measure the impact of the LLMs on the final performance, not compensated and leveled by complex agent architectures.
@@
The agent processed a given user task without Human-In-The-Loop involvement. Each task normally requires series of tool calls to solve and the agent does not always find the shortest solutions, so we allowed up to 25 iterations “LLM suggests tool call(s) – LLM receives the tool(s) response”. Less iterations were not enough for some complex tasks, while allowing more steps did not increase success rates. Experiments indicated that an increase in temperature leads to the agent being stuck looping over the same steps and to lower task completion rates, so, all LLMs with enabled temperature parameter were tested at temp. 0.
@@
The agent system prompt sets the persona, briefly describes the setting, and the rules for solving the tasks (complete prompt in Appendix). The agent was instructed not to ask the user for clarifications and not to provide code to solve the task, but to respond with the resulting map, the names of the data sets used and explanations. The set of rules instructs when to reject the task and when to leave the solution empty (meaning that no tools are needed to answer user’s question), when to consider the data available (if the available data do not fully cover the geography asked in the task), and how to select columns for mapping data if the user did not specify (implied conditions frequent in production user questions).
@@
The text of the task is supplied as the next message for the LLM Chat API as the user prompt.
@@
Agent is provided with 23 tools (specifications are in the Appendix) that can read tabular data, geospatial data in vector or raster formats, perform operations on dataframes and geodataframes (merge, filter by categorical or numerical values), their columns, perform spatial operations on geodataframes and rasters (spatial joins, filtering by locations, creating buffers, generate contour lines from raster), visualize geodataframes, generate choropleth maps and heatmaps. The tools leverage GIS-specific libraries like GeoPandas den Bossche et al. (2024), Rasterio Gillies et al. (2013–), GDAL Rouault et al. (2025), Shapely Gillies et al. (2025), and common data analysis libraries like Matplotlib, Plotly, Pandas, and Numpy.
To distinguish the situations in which the agent stops solving the task because it decides that the task could not be solved with the provided data sets / tools, the agent was equipped with an additional tool rejecttask that simply returned the message that the task is unsolvable.
@@
When a tool result included a dataframe or geodataframe, information about its length, columns, data types, and non-empty cells was included in the tool response message. Where dataframes or geodataframes needed to be passed between tools, they were stored in the data-store dictionary accessible to all tools.
@@
To make the experimentation consistent, we opted to store the snapshots of datasets in the local experimentation environment rather than fetching data from online APIs on every run. This allowed us to avoid variability of benchmarking due to possible data API connection issues, data API changes, and data changes behind the scenes.
Snapshots of data sets were organized as dictionaries with short descriptive titles as keys and filenames as values, one dictionary per type of data. When the tool designed to load the respective data was called, LLM had to choose the tool argument, the data set to load, from the list of the keys of the appropriate dictionary.
@@
The data included: Statistical data: panel data in csv format (18 datasets): World Bank data on economic and population indicators, tuberculosis cases at subnational levels. Geometry data: vector data in shapefile format (21 datasets): polygon geometries (national or subnational units, lakes), points geometries (wildfire locations, urban places, Amtrak stations, earthquake occupancies, seaports, mineral extraction facilities, power stations), line geometries (rivers, railways). Raster data: (11 rasters): accumulated snow cover, flood extents, population data.
4 Benchmark set
To represent typical use by GIS professionals of a GIS assistant, tasks were single-sentence questions in natural language that might be asked by someone who needs a map, a number, a visualization for drafting an analytical research note, creating an action plan in international development, cross-checking information, etc. Some tasks were naturally ambiguous or missing information.
@@ 4 Benchmark set
We created a set of 200+ tasks, grouped by type and complexity. Based on Wald’s method, this sample size corresponds approximately to the 5% margin of error for estimating a binary match rate with the 90% confidence. Each group included tasks that could be solved with the tools/datasets provided (solvable tasks), tasks that looked similar to solvable but could not be solved with the tools or datasets provided (unsolvable tasks). The groups of tasks are as follows:
Merge-visualize
@@
(36 tasks): require to join tabular numeric data (panel data for one or multiple indicators, e.g. countries-by-years) with a geographic geometry data and make a choropleth or bivariate map. Example: Map the relationship between GDP per capita and electric power consumption per capita globally”. This is the most straightforward group, the tasks mostly require selection of correct datasets, and two (three in case of bivariate map) tool calls. Only 6 of these tasks were not solvable with given tools/data. Due to simplicity of the tasks, their content and solutions have limited variation; therefore, a limited number of them needed to have enough coverage.
Process-merge-visualize
@@
(56 tasks): require some processing of either tabular or geometry data (filtering or operations on columns) and then creating a choropleth or bivariate map. Example: Compare rural population percentages in Western Asian countries”. The tasks in this group are more complicated, require the use of knowledge of the world geography and the use of a wider variety of tools. Half of the tasks, 28, are not solvable with given tools/data.
Spatial operations
@@
(53 tasks): require spatial joins, creating buffers around selected geometries, checking overlaps between rasters, performing calculations on rasters, finding values from raster using geometry data or filtering geometries based on raster values, calculating distances. They might require a numerical answer, and the creation of actual maps might be optional. This group is more challenging than the previous two, as it requires the agent to work with different types of data set in one task, navigate the full set of available tools, and maintain the overall logic of the solution through all its steps. Example: How many people live within 1 km from a railway in Bangladesh?”. In this group, 23 of the tasks were not solvable.
Heatmaps, contour lines
@@
(54 tasks): required performing spatial operations (selection, overlaps, etc.), understanding the contents of rasters, and making a heatmap or generating contour lines as one of the steps. Example: Create a heatmap of earthquake occurrences in regions with high population growth”. This was the most challenging set, as the agent had to figure out if the heatmaps and contour lines tools were actually able to solve the task and to make correct decisions on a complex set of arguments. In this group, 41 tasks ended up being unsolvable with available data/tools, which was a test for LLMs ability to recognize an unsolvable task and generate a rejection rather than a hallucinated answer.
4.1 Tasks creation
@@
For each group, seed tasks were sourced from industry professionals and publications. Then more tasks were generated by using the Sonnet 3.5 model based on the seed tasks, the description of the tools, and the lists of available datasets. Then tasks were generated that should look similar to solvable tasks but could not be solved using the provided data and tools.
All the generated tasks were then manually reviewed to keep only high-quality tasks in the set.
4.2 Tasks annotation
@@
Each of the 200+ tasks was annotated with one or multiple ground-truth reference solutions. To create the annotation, we have started by generating the draft reference solutions by the agent using GPT-4o or Sonnet 3.5 models. We then manually reviewed and rewrote every solution.
@@
One reference solution could be (1) Sequence of tool calls with arguments; similar to a Python code snippet (2) Empty sequence [] of tool calls meaning that the task can be solved from the LLM’s world knowledge (3) A single special tool call reject_task() meaning that the task should be rejected as it cannot be solved with given tools/data.
A nonempty reference solution looked as follows for the evaluator:
@@
#toolcall1 comment
toolcall1(arg11=..., arg12=..., ...)
#toolcall2 comment
toolcall2(arg21=..., arg12=..., ...)
Approximately 75% of the solvable tasks have reference solutions consisting of 4-6 steps (tool-calls) and  25% consist of more steps (up to 14 steps).
We used comments in annotations as a way to record information important for the evaluator agent to make a judgment about whether the candidate solution and the reference solution are semantically equivalent. For example, telling which variations in the solution or step are irrelevant or important.
@@
One task sometimes allowed multiple solutions because tasks in practice are often ambiguous and multiple analytical approaches are possible for them. For example, task Compare the freshwater withdrawal between African countries with and without significant railway networks” can be solved: (1) by rejecting because it is not specific enough (user need to clarify what’s “signigicant” railway networks; (2) by making separate maps of railway networks and water withdrawals for African countries; (3) by combining the data and mapping bivariate map to show the relation; (4)by implying some threshold and computing statistical metrics for 2 groups of countries.
5 Evaluation approach
5.1 Matching-based evaluator

### v2

# GeoBenchX: Benchmarking LLMs in Agent Solving Multistep Geospatial Tasks

@@ Abstract.
This paper establishes a benchmark for evaluating tool-calling capabilities of large language models (LLMs) on multi-step geospatial tasks relevant to commercial GIS practitioners. We assess eight commercial LLMs (Claude Sonnet 3.5 and 4, Claude Haiku 3.5, Gemini 2.0 Flash, Gemini 2.5 Pro Preview, GPT-4o, GPT-4.1 and o4-mini) using a simple tool-calling agent equipped with 23 geospatial functions. Our benchmark comprises tasks in four categories of increasing complexity, with both solvable and intentionally unsolvable tasks to test rejection accuracy. We develop a LLM-as-Judge evaluation framework to compare agent solutions against reference solutions. Results show o4-mini and Claude 3.5 Sonnet achieve the best overall performance, OpenAI’s GPT-4.1, GPT-4o and Google’s Gemini 2.5 Pro Preview do not fall far behind, but the last two are more efficient in identifying unsolvable tasks. Claude Sonnet 4, due its preference to provide any solution rather than reject a task, proved to be less accurate. We observe significant differences in token usage, with Anthropic models consuming more tokens than competitors. Common errors include misunderstanding geometrical relationships, relying on outdated knowledge, and inefficient data manipulation. The resulting benchmark set, evaluation framework, and data generation pipeline are released as open-source resources  [^1] Available at https://github.com/Solirinai/GeoBenchX, providing one more standardized method for the ongoing evaluation of LLMs for GeoAI.
Large Language Models (LLMs), Geospatial Analysis, Function Calling, LLM-as-Judge Evaluation, GIS Tools, Benchmarks, Spatial Reasoning, Tool-based Agents
journalyear: 2025copyright: acmlicensedconference: The 1st ACM SIGSPATIAL International Workshop on Generative and Agentic AI for Multi-Modality Space-Time Intelligence; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 1st ACM SIGSPATIAL International Workshop on Generative and Agentic AI for Multi-Modality Space-Time Intelligence (GeoGenAgent ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3764915.3770721isbn: /2025/11ccs: Computing methodologies Artificial intelligenceccs: General and reference General conference proceedingsccs: General and reference Evaluationccs: Computing methodologies Intelligent agentsccs: Computing methodologies Multi-agent systemsccs: Computing methodologies Information extraction
1. Introduction
@@
In this paper, we create a benchmark set and describe a simple LLM-as-Judge based methodology to compare the tool-calling performance of several leading commercial large language models (LLMs) with function-calling abilities for solving a range of multi-step geospatial tasks. Geospatial tasks are distinct in that they require LLMs to have good factual geographical knowledge, understanding how to work with spatially defined data, what are coordinate reference systems, and various types of spatial data. We used a simple tool-calling agent to specifically assess the LLMs capabilities and have tested several models: Anthropic’s Claude Sonnet 3.5, 4, Claude Haiku 3.5; Google’s Gemini 2.0 Flash, Gemini Pro 2.5 Preview; OpenAI’s GPT-4o, GPT-4.1, o3-mini and o4-mini.
@@
Our focus is on the commercial GIS practitioner requirements, which influences the choice of tasks, models and evaluation framework. Unlike other adjacent benchmarks, we focus on tasks needed for socio-economic analysis and mapping for reports, policy papers, news articles, and the like: tasking agent to make various types of maps, to find geographical objects by some conditions, to count objects within certain geographic boundaries. For these tasks, practitioners would use LLM equipped with GIS tools in a chat environment. Also, importantly for practitioners, we test the ability of LLMs to correctly recognize when the task cannot be solved with given tools and inform the user to reduce the risk of hallucinated outputs. Next, we selected the most current advanced commercial LLMs used in the industry setting for benchmarking. Lastly, the evaluation method we use allows for automation of the ongoing evaluation which is increasingly important with frequent models updates in commercial settings.
@@
The paper is structured as follows. First, we describe the task-solving agent and the set of tools. Second, we describe how the benchmark set of tasks was created and its composition. Third, we detail the approach to evaluation of the results of the task-solver agent using a panel of LLM-as-Judge evaluators and testing the evaluator performance itself. Finally, we compare the performance metrics of the solver agent based on different LLMs and the cost efficiency in terms of token usage.
2. Related works
@@
Research in the field of GeoAI, a part of spatial data science at the intersection with Artificial Intelligence (AI) (Janowicz et al., 2020), got a boost with recent advances in generative AI and the deployment of accessible LLMs. GeoAI researchers actively explore the capabilities of the available models in relation to geospatial data science, design their applications, and methods to improve their performance. The avenues explored include analysis of the ability of LLM (GPT-4) for spatial reasoning in travel networks, distance and elevation estimations, route and supply chain analysis (Roberts et al., 2023); integration of geography knowledge in pre-trained language models (Ramrakhiyani et al., 2025), creating autonomous LLM-based GIS applications (LLM-Geo) (Li and and, 2023), GIS Copilot for GIS software (Akinboyewa et al., 2024), integrating LLM and cartography tools into MapGPT (He et al., 2024), testing LLMs in styling maps using reference visuals (Wang et al., 2025), combining several LLMs to increase the efficiency of spatial data processing in the multi-agent framework for automated processing of vector data in shapefile format (Lin et al., 2024), benefits of generative AI agents in smart mobility(Xu et al., 2024), fine-tuning LLM to process natural language queries into executable code(Mansourian and Oucheikh, 2024). There is a benchmarking of LLM performance in spatial tasks in progress, with analysis of LLMs’ geospatial knowledge(Bhandari et al., 2023; Renshaw et al., 2025), of GIS coding abilities(Gramacki et al., 2024), the recent study testing models on several groups of tasks that require geographic, GIS, and spatial knowledge, including changes in their performance in response to changes in prompts (Xu et al., 2025).
3. Experiment setup
3.1. Agent
@@
The experiment setup is based on the pre-built Langgraph ReAct agent (LangChain AI, 2024) equipped to call one of the selected LLMs with a list of tools (functions) to solve tasks. We intentionally choose a simple agent architecture to measure the contribution of the LLMs themselves into the final performance, not compensated and leveled by complex agent architectures. Typical practical requirements are that the variability in outputs is undesirable, but consistency in responses is critical, so the models were tested at temperature 0, which reduces variability and increases consistency of outputs (OpenAI, 2025; Google Cloud, 2025).
@@
The agent processed a given user task without Human-In-The-Loop involvement. Each task normally requires series of tool calls to solve and the agent does not always find the shortest solutions, so we allowed up to 25 iterations of “LLM suggests a tool call –¿ LLM receives the tool response”. This number was selected after initial trials of the agent that showed that fewer iterations were not enough for some complex tasks, while allowing more steps did not increase the success rates (accuracy).
@@
The agent is supplied with a simple system prompt that sets the persona, briefly describes the setting, and the rules for solving the tasks. The agent is explicitly instructed not to ask the user for clarifications and not to provide code to solve the task, but to reply with the resulting map, names of the datasets used, and explanations. The set of rules contains instructions on when to reject the task and when to leave the solution empty (meaning that no tools are needed to answer user’s question), when to consider the data unavailable (if the available data do not fully cover the geography asked in the task), and how to select columns for mapping data if the user did not specify (implied conditions frequent in production user questions).
@@
The text of the task is supplied as the next message for the LLM’s Chat API as the user prompt.
3.2. Agent tools
@@
To enable the agent to solve the tasks, it is provided with 23 tools to implement the necessary data manipulations, calculations, and visualizations or to report a task unsolvable. Tools can read tabular data, geospatial data in vector or raster formats, perform operations on dataframes and geodataframes (merge, filter by categorical or numerical values), their columns, perform spatial operations on geodataframes and rasters (spatial joins, filtering by locations, creating buffers, generate contour lines from raster), visualize geodataframes, generate choropleth maps and heatmaps. The tools leverage GIS-specific libraries like GeoPandas (den Bossche et al., 2024), Rasterio (Gillies et al., 13), GDAL (Rouault et al., 2025), Shapely (Gillies et al., 2025), and common data analysis libraries like Matplotlib, Plotly, Pandas, and Numpy. To distinguish the situations in which the agent stops solving the task because it thinks that the task could not be solved with the datasets / tools provided, the agent was equipped with an additional tool rejecttask that simply returned the message that the task is unsolvable.
@@
The outputs of the tool functions were formatted as natural part of the tool response message. When a tool result included a dataframe or geodataframe, information about its length, columns, data types, and non-empty cells was included in the tool response message. Where dataframes or geodataframes needed to be passed between tools, they were stored in the data-store dictionary accessible to all tools.
@@
To avoid variability of benchmarking due to possible data API connection issues, data API changes, and data changes behind the scenes, we opt to store the snapshots of datasets in the local experimentation environment rather than fetching data from online APIs every time.
For each task, the conversation history between LLM and tools along with any plotted outputs is saved as html file making the final LLM response, maps and intermediate outputs available for debugging.
3.3. Datasets
The datasets were used as-is after downloading from online APIs, except for two modifications made to ease development: country names in the vector geodataset were harmonized with statistical datasets for easier visual identification of filtering issues, and the NoData value in snow cover rasters was unified. These changes were not required for the benchmarking itself.
The data included:
Statistical data: panel data in csv format (18 datasets): data on economic and population indicators, data on CO and greenhouse gas emissions, tuberculosis cases at subnational levels.
@@ 3.3. Datasets
Geometry data: vector data in shapefile format (21 datasets): polygon geometries (national or subnational units, lakes), points geometries (wildfire locations, urban places, Amtrak stations, earthquake occurrences, seaports, mineral extraction facilities, power stations at global or national level), line geometries (rivers, railways continent or national extent).
Raster data: (11 rasters): accumulated snow cover, flood extents, population data at national level.
4. Benchmark set
@@
We created a set of 200+ tasks, grouped by type and complexity. For estimating a binary match rate at the level of match we see in results section (6) – this sample size corresponds approximately to 5% margin of error with 90% confidence according to Wald method. Each group included tasks that could be solved with the tools and datasets provided (solvable tasks), tasks that looked similar but could not be solved with the tools or datasets provided (unsolvable tasks). For debugging purposes, we also added a few tasks that did not require use of the provided tools or datasets to solve.
4.1. Groups
Depending on the complexity of data processing or geoprocessing, the task groups are as follows:
@@ 4.1. Groups
Merge - visualize (36 tasks): require to join tabular numeric data (panel data for one or multiple indicators, e.g. countries-by-years) with a geographic geometry data and make a choropleth or bivariate map. Example: Map the relationship between GDP per capita and electric power consumption per capita globally”. This is the most straightforward group, the tasks mostly require two-three tool calls. Only 6 of these tasks were not solvable with given tools/data. Due to simplicity of the tasks, there is limited variation in their content and solutions and a smaller number provided enough coverage.
@@
Process - merge - visualize (56 tasks): require some processing of either tabular or geometry data (filtering or operations on columns) and then creating a choropleth or bivariate map. Example: Compare rural population percentages in Western Asian countries”. The tasks in this group are more complicated, require the use of knowledge of the world geography and the use of a wider variety of tools. About half of the tasks, 24, are not solvable with given tools/data.
@@
Spatial operations (53 tasks): require spatial joins, creating buffers around selected geometries, checking overlaps between rasters, performing calculations on rasters, finding values from raster using geometry data or filtering geometries based on raster values, calculating distances. They might require a numerical answer, and the creation of actual maps might be optional. This group is more challenging than the previous two, as it requires the agent to work with different types of datasets in one task, navigate the full set of available tools, and maintain the overall logic of the solution through all its steps. Example: How many people live within 1 km from a railway in Bangladesh?”. In this group, 22 of the tasks were not solvable.
@@
Heatmaps, contour lines (54 tasks): required performing spatial operations (selection, overlaps, etc.), understanding the contents of rasters, and making a heatmap or generating contour lines as one of the steps. Example: Create a heatmap of earthquake occurrences in regions with high population growth”. This was the most challenging set, as the agent had to figure out if the heatmaps and contour lines tools were actually able to solve the task and to make correct decisions on a complex set of arguments. In this group, 27 tasks were not solvable.
4.2. Tasks creation
@@
For each group, the seed tasks were sourced from industry professionals and publications. Then more tasks were generated by using the Claude 3.5 Sonnet model based on: the seed tasks, description of the tools, and the lists of available datasets. Unsolvable and solvable tasks were generated separately. The generated tasks were then manually reviewed to keep only high-quality tasks in the set.
4.3. Tasks annotation
@@
To allow the evaluation of LLM-generated solutions, each task was annotated with the ground truth, reference solution, formatted as snippets of Python code (see more in Section 5). To create the annotation, we have started by generating draft reference solutions by the agent using GPT‑4o or Claude 3.5 Sonnet models. We then manually reviewed and rewrote every solution. Each of the 200+ tasks was annotated with one or multiple ground-truth reference solutions.
Reference solution could be
  1. (1)
    Sequence of tool-calls with arguments; similar to a python code snippet
  2. (2)
    Empty sequence [] of tool-calls meaning that the task can be solved from the LLM’s world knowledge
  3. (3)
@@
    A single special tool call reject_task meaning that the task should be rejected as it cannot be solved with given tools/data
A non-empty reference solution looked as follows for the evaluator:
@@
toolcall1(arg11=..., arg12=..., ...) #call1 comment
toolcall2(arg21=..., arg12=..., ...) #call2 comment
@@
One task sometimes allows multiple solutions because tasks in practice are often ambiguous and multiple analytical approaches are possible for them based on the provided data and tools. For example, task Compare the freshwater withdrawal between African countries with and without significant railway networks” can be solved:
  • by rejecting because it’s not specific enough (user needs to clarify what’s “significant” railway networks);
  • by making separate maps of railway networks and water withdrawals for African countries;
  • by combining the data and mapping bivariate map to show the relation;
  • by implying some threshold and computing statistical metrics for 2 groups of countries.
We also used comments as a way to record some information important for the evaluator to make a judgment about whether the candidate solution and the reference solution are semantically equivalent. For example, providing alternative arguments to the tool call, specifying variations in arguments or marking optional steps.
Approximately 74% of the solvable tasks have reference solutions consisting of 4-6 steps (tool-calls) and  28% consist of more steps (up to 14 steps). The numbers do not total to 100% as a task can have multiple reference solutions.
5. Evaluation approach
5.1. Matching-based evaluator